OpenStudy (anonymous):

Marlon has just won some money on a game show! He has the option to take a lump sum payment of $625,000 now or get paid an annuity of $4,500 per month for the next 15 years. Assuming the growth rate of the economy is 3.9% compounding annually over the next 15 years, which is the better deal for Marlon and by how much?

OpenStudy (anonymous):

I think it's going to be the annuity but I've been having issues finding the Future Value of an Annuity

OpenStudy (tkhunny):

What's stopping you from calculating the Present Value of the Annuity Payments and simply comparing the Calculated PV with the $625,000?

OpenStudy (anonymous):

Do you know the future value of an annuity formula?

OpenStudy (tkhunny):

Why do you care about a Future Value? $625,000 is NOW, not in the future. i = 0.039 -- This is given in the problem statement. j = i/12 = 0.00325 -- The 12 (monthly) is given in the problem statement. v = 1/(1+j) -- Monthly Interest Discount Factor (or Inflation Factor) PV of All Monthly Payments = \(4500 + 4500v + 4500v^{2} + ... + 4500v^{180-1}\) $4,500 monthly is given in the problem statement. 180 months is given int eh problem statement as "15 years". Add them all up! PV of All Monthly Payments = \(4500(1 + v + v^{2} + ... + v^{180-1})\) PV of All Monthly Payments = \(4500\cdot \dfrac{1-v^{180}}{1-v}\) Can you see all that and can you then calculate the value?